# Tuning without Peeking: Provable Privacy and Generalization Bounds for LLM Post-Training

### v1

@@ 5.2 Discussion of experimental results
Robustness and statistical validation. Our experiments show that BBoxER achieves statistically significant performance gains without overfitting across a range of BBO algorithms (see P-value validation in App. J), even on Instruct models. Consistent with theory, the Bet-and-Run 3-OneFifth further improves performance and allows for larger budgets. More discussion in App. I.
@@
Comparison. The goal of our experiments is not to outperform state-of-the-art benchmarks, but to establish that effective learning occurs in BBoxER—thereby validating its applicability and enabling the theoretical guarantees to take effect. That said, it is still informative to situate our results within the broader landscape of current methods for reasoning. Recent approaches such as GRPO [45] have proven highly effective on reasoning tasks, though few explicitly target small-scale datasets like GSM8k or explicitly study out of domain transfer like to GSM+. A notable exception is [85], which gives, for few-shot evaluation, results comparable to ours on Llama3.1-8B-Instruct (see App. I).
Computational cost. Our approach is lightweight: it relies only on the forward passes of LLMs without computing and storing any gradients or large optimizer state. This allows us to fit a large batch size during training. The overhead of the internal state of the BBO algorithm is negligible compared to the time required for a forward pass on the training dataset.
@@ 6 Conclusion
This paper introduced BBoxER, a principled post-training method that comes with strong bounds against overfitting, guarantees privacy by design, ensures resilience to data poisoning, robustness to data extraction, without any assumption on the downstream algorithm. Thanks to the use of comparison-based BBO algorithms for retrofitting, the training set is compressed into a small number of bits, yielding strong theoretical guarantees on overfitting, robustness, and privacy. Despite this compression, BBoxER empirically demonstrates effective learning and transfer capabilities on GSM8k with contemporary LLMs. Theoretical insights are matched by experimental performance, especially with Bet-and-Run algorithm combinations under full attention broadcast. Crucially, this work is the first to bridge the LLM training landscape with the full BBO zoology. This connection, grounded in formal results, paves the way for safer, robust, and more adaptable language models, and opens the door to broader adoption of privacy-preserving training at scale, including completely remote A/B testing and training models directly on user satisfaction without any signal about individual usage.
Limitations. Our theoretical bounds are not always tight, leaving room for refinement. By design, BBoxER is ill-suited for data retrieval tasks due to its compressed optimization trace. Finally, while challenging, scaling BBoxER to larger parameter spaces appears achievable with further work.
@@ Appendix A Extended Related Work
Data poisoning alters training data to inject specific behaviors, either to degrade performance or to create targeted backdoors [19]. This includes subtle manipulations—e.g., poisoning just $0.001\%$ of training data in medical applications [90]—that remain undetected by conventional metrics. Attacks have proven transferable across quantized and full-precision models [62]. Recent work shows that inserting only a few percent of malicious preference pairs in RLHF data can dramatically shift model behavior [15, 14], such as increasing output verbosity or sentiment polarity. These changes are activated by specific triggers and survive quantization. Theoretical work [80, 63] confirms that even small adversarial corruptions can be highly effective, especially in overparameterized models. However, certified robustness against poisoning in LLMs remains an open problem. Compared to e.g., [63], we provide an explicit algorithm with the desired stability assumptions.
@@
LLMs are exposed to risks of privacy and data extraction attacks [4]. DP-SGD [65] introduces formal guarantees by adding Gaussian noise to clipped gradients, achieving $(\varepsilon,\delta)$-DP under strict accounting [91]. Nevertheless, it suffers from increased compute, memory use, and degraded utility. Extensions like adaptive clipping and layer-wise noise scaling aim to mitigate this [91]. PATE [66] aggregates teacher votes to train a student model privately. Parameter-efficient strategies like DP-BiTFiT [92], soft prompts [93], and selective DP [94] offer better trade-offs. Our method avoids the need for noisy gradient updates by enforcing a compression bottleneck and comparison-based optimization, achieving $(\varepsilon=0,\delta=0)$ for all critical content, and targeting high-fidelity memorization pathways directly. This aligns with broader insights that data retention precision is a root cause of privacy and misuse [95, 35]. Yu et al. [96] proposed a meta-framework for differentially private fine-tuning pretrained language models, achieving state-of-the-art privacy-utility tradeoffs on standard NLP tasks. Their approach was inspired by the success of parameter-efficient fine-tuning methods, allowing for simpler and sparser algorithms while maintaining high utility and providing rigorous privacy guarantees. In this case, by leveraging DP-SGD with careful gradient clipping and noise addition, the framework ensures that the influence of any single data point on the model’s parameters is bounded, thus protecting individual privacy. However, although the methods achieve strong performance, there is still a noticeable drop in accuracy compared to non-private fine-tuning, especially for smaller models or more complex tasks. Furthermore, the effectiveness of their DP fine-tuning approach is sensitive to hyperparameters such as learning rate, clipping norm, and noise multiplier, illustrating that careful fine-tuning is required to balance privacy and utility. Other works have focused on privacy-preserving strategies for instruction tuning and alignment of large language models. For instance, Yu et al. [97] introduced a two-stage framework that addresses privacy risks not only during training, but also in the instruction collection phase. In this case, they first fine-tune a pretrained LLM using differentially private optimization (DP-Adam) to generate synthetic user instructions, and then apply a DP histogram-based distribution matching technique to resample instructions such that they resemble the distribution of the original private data. This approach ensures that both annotators and the final alignment pipeline never see raw user inputs, thereby preventing both direct exposure and downstream memorization. Applied to LLaMA [98] and Phi-models [99], their approach achieved state-of-the-art performance in both supervised fine-tuning and RLHF settings, with utility comparable to non-private baselines. In particular, they highlight the importance of privacy-preserving data preparation in addition to model training.
@@
Reasoning tasks like GSM8k [52], SVAMP [84], and MATH [53] highlight the limitations of standard LLM (pre)training. Post-training approaches for reasoning tasks are now commonplace in LLM training. Enhancements via chain-of-thought prompting [100], complex reasoning extensions [101], and in-context learning [102, 103] are effective but often brittle. Reinforcement learning approaches such as GRPO [45] allow fine-tuning using unsupervised signals, opening the door for real-time user-driven improvement [104]. LoRA [105] and robust tuning [106] enhance adaptability but are prone to hallucination [107] and exploitative behaviors [108, 35]. Our approach combines low-rank adaptations with retrofitting techniques and avoids reward-based fine-tuning. It limits overfitting risks and prevents the amplification of memorized or poisoned content by emphasizing rank-based signal aggregation over token-level supervision. A somewhat complementary approach is self-improving [109]. In such a case, a large language model tries to improve its performance by self-fine-tuning with synthetic data generated by itself. Nevertheless, such approaches have mostly focused on performance, and they are not risk-free from data poisoning or leaking private data.
@@ Appendix B Examples of Modifiers
The BBoxER framework is described in Alg. 1. This algorithm has various hyperparameters, such as a budget, a black-box optimization algorithm (see Sec. 3.2), and a $modified(m_{0},x)$ function which specifies how a vector $x$ proposed by the black-box optimization method is used to modify an initial model $m_{0}$. Here we present a few examples of such $modified(\dots)$ functions, involving broadcast, low-rank, or full updates.
@@ B.1 Examples of Modifiers for One Layer
In the present section $\times$ refers to pointwise multiplication and @@@@ refers to matrix multiplication.
@@
2:  $model=copy(m_{0})$
3:  $model.output\_matrix=model.output\_matrix\times\exp(0.01\times x)$
4:  Return $model$
@@
Other update formulas include $+0.01\times x$, ×$\times(Id+0.01\times x)$, and other constants than $0.01$. However, in most cases, we will use vectors $x$ rather than matrices, as explained in low rank and broadcast.
@@
When updating entire weight matrices, we prefer to work with low-rank (LoRA) updates (rank 1 in fact), defined as follows, when $model.output\_matrix$ has shape $(hidden\_dim,vocab\_size)$ and $x$ has dimension $hidden\_dim+vocab\_size$:
@@
2:  $model=copy(m_{0})$
3:  $x_{1},x_{2}=split(x)${$x_{1}$ has shape $(hidden\_dim,1)$ and $x_{2}$ has shape $(1,vocab\_size)$}
4:  $model.output\_matrix=model.output\_matrix\times\exp(0.01\times x_{1}@x_{2})$
5:  Return $model$
@@
In some cases, we use broadcast rather than low-rank updates. $modified(m_{0},x)$ is then defined as follows, where $x$ as a row vector is repeated $M$ times to fit the shape of the matrix we want to update. In Sec. 5 $x$ has dimension $hidden\_dim$ and $M=hidden\_dim$:
@@
2:  $model=copy(m_{0})$
3:  $x_{1}=(1,\dots,1)$
4:  $model.output\_matrix=model.output\_matrix\times\exp(0.01\times x_{1}^{t}@x)$
5:  Return $model$
In App. I, $x$ has dimension $vocab\_size$ when broadcast updates are applied to the output layer.
@@ B.2 Modifying Several Layers
In the examples above, we consider updates of one layer only. However, this can be naturally extended to updates of $l$ layers. Then, $x$ is split into $x_{1},\dots,x_{l}$, and each $x_{i}$ is used for updating a specific layer, as in Sec. B.1.
In our experiments in Sec. 5 we performed the following updates:
@@
For all these experiments the update constant is $0.01$. A few additional experiments in App. I use a different constant.
@@ C.2 Probably Approximately Correct (PAC) Learning: A Brief Overview
  • Sec. C.2.1 presents the Bonferroni correction: whereas “non-uniform” large deviation inequalities estimate the difference between an empirical error and a generalization error for a single predefined classifier (for example when testing a single model on a test set), the Bonferroni correction provides bounds applicable uniformly on a finite set of classifiers (e.g., for selecting, with a validation set, a classifier in a list of classifiers obtained on the training set).
  • Sec. C.2.2 generalizes the Bonferroni correction to infinite families of classifiers: VC-dimension bounds, and others, are the most well-known bounds. They are based on limiting the “capacity” of a space of function, even when it is infinite.
  • Sec. C.2.3 presents other bounds for infinite families of functions, taking into account the optimization algorithm used for selecting a classifier: the idea is that even if the space of functions is large, we can get bounds if the optimization algorithm satisfies some stability assumptions. This last part of the state-of-the-art is the closest to our work: we get non-trivial bounds on deep learning models, independently of the possibly huge number of parameters.
@@ C.2.1 Bounds on the Generalization Loss based on Bonferroni
Let us see the generalization of large deviation inequalities seen in Sec. C.1.
@@
Thm 1 extends this to more sophisticated cases, with $N(\omega,a,b)=\pi_{i}k_{i}(\omega,a,b)>>b$ (in many cases, exponential in $b$ e.g. $N(\omega,a,b)\leq 2^{b}$ and $k_{i}\leq 2$) instead of $N(\omega,a,b)=b$. In spite of being larger, this bound, exponential in $b$, is tighter than bounds based on capacity.
@@ C.2.2 Bounds on the Generalization Loss based on Capacity: Statistical Learning Theory from the 90s
(assuming that $\hat{x}$ minimizes $\widehat{L}(\widehat{x})$). When using shattering coefficients instead of VC-dimension, one can read:
@@ Appendix D Black-box Optimization Algorithms
The BBoxER algorithmic framework, as illustrated in Figure 2, embeds an algorithm from the family of block-box optimization algorithms, which we now describe in detail.
Dataset $\mathcal{D}$RetrofittingcompressorCompression $c$(hundreds orthousands of bits)RetrofittingmodifierFinal modelInitial model
Figure 2: Schematic view of BBoxER Alg. 1. Our approach compresses the data relative to the current model, leading to a list $c$ of $b$ bits (Eq. 3, $b$ is the total budget). The final model depends only on those $b$ bits, giving perfect control over the number of bits transferred from the dataset to the model. In BBoxER, budgets $b$ are typically between 60 and 1500 bits, i.e., very small compared to the dataset size. The “Retrofitting compressor” refers to the part of Alg. 1 which outputs the list $c$ of outputs of the $Compare(\dots)$ function: $c$ represents a compression of the dataset $D$ conditional on the initial model. “Retrofitting modifier” is the part of Alg. 1 which outputs the final model, given the list $c$ and the initial model.
@@ D.1 Evolutionary Algorithms
For the reasons given in Sec. 3.2, only Evolutionary Algorithms (EAs) have been used in the present work, as they are (at least for those uses in the present document) comparison-based, and hence de facto comply with the finite branching factor condition, which is mandatory for the theoretical results of Sec. 4 to hold. Sec. D.2 discusses this choice, and its possible relaxation, while the current section introduces EAs more thoroughly, and details all the comparison-based algorithms used in the experiments of Sec. 5.
@@
The first adaptive ES uses the so-called “one-fifth success rule” [119, 117], established by theoretical analyses on simple linear and quadratic objective functions, on which the optimal behavior is reached for a success rate (offspring better than parent) of approximately $0.2$. This is available in Nevergrad as the (1+1)-ES with 1/5th rule, named in this paper OneFifth, whose pseudo-code is given in Alg. 2-left.
@@
When handling a full population ($\mu>1$), both CMA-ES and D-CMA apply the adaptive Gaussian mutation to a linear recombination of the $\mu$ individuals. The historical CMA-ES used the mean (all weights of the linear recombination are $1/\mu$), whereas the state-of-the-art CMA-ES uses weights that decrease with the rank of each individual. In the latter case, more information is used, and the branching factor increases (see Sec. E.5).
As noted in Sec. 3.2, the simplest ESs (or EAs) are the (1+1)-EAs: the single current individual is modified, and the best of both becomes the new current one. While the already-mentioned OneFifth was the first historical successful adaptive EA (see again Alg. 2-Left), another family of (1+1)-EAs is directly derived from the historical GAs (Genetic Algorithms), that optimize bitstrings: the mutation typically flips each bit with a given probability. In the continuous case, the bitflip is turned into a “Gaussian flip”: each value is replaced by a value sampled from a standard normal mutation (${\cal N}(0,1)$ in Alg. 2). Several variants of such “discrete-inspired” algorithms exist in Nevergrad, depending on the probability of applying this Gaussian flip to each variable. The Discrete algorithm exactly mimics the standard setting in GAs by adopting a fixed $1/d$ probability; the Lengler [121, 122] algorithm uses a gradually decreasing schedule. The Uniform-Disc [123] algorithm uses a probability that is uniformly sampled in ]0,1]]0,1]] 0 , 1 ], and the FastGA [124] uses an $\alpha/d$ probability where $\alpha$ is sampled from a power-law distribution.
@@
Adaptive One-fifth $(1+1)$-ES 0: initial point $x_{0}\in\mathbb{R}^{d}$; budget $b\in{\mathbb{N}}$ 0: step-size $\sigma>0$ 1:$x=x_{0}$ 2:for $i\in\{1,\dots b\}$ do 3:    $x^{\prime}=x+\sigma{\cal N}(0,1)$ 4:    if $x^{\prime}$ better than $x$ then 5:       $x=x^{\prime}$ 6:       $\sigma=2\sigma$ 7:    else 8:       $\sigma=2^{-\frac{1}{4}}\sigma$ 9:    end if 10:end for Generic Discrete-inspired $(1+1)$-EA 0: initial point $x_{0}\in\mathbb{R}^{d}$; budget $b\in{\mathbb{N}}$ 1:$x=x_{0}$ 2:for $i\in\{1,\dots b\}$ do 3:    $x^{\prime}=x$ 4:    while $x^{\prime}=x$ do 5:       $x^{\prime}=x$ 6:       for $j\in\{1,\dots,d\}$ do 7:          With probability $p(i,b,d)$$x^{\prime}_{j}={\cal N}(0,1)$ 8:       end for 9:    end while 10:    if $x^{\prime}$ better than $x$ then 11:       $x=x^{\prime}$ 12:    end if 13:end for
Algorithm 2 The two typical (1+1)-EAs.
Table 3: Bound on the average branching factor for various algorithms (see Sec. E.4, Sec. E.5, Sec. E.6).
@@
OneFifth $2$
Lengler $2$
FastGA $2$
Portfolio $2$
@@
DE, PSO $2$
DE, PSO with curr-to-best $3$
CMA-ES/D-CMA $2$
@@
These algorithms were chosen because of their finite branching factor: Tab. 3 lists the branching factors of all cited algorithms. As noted in Sec. 3.2, the branching factor is 2 for all (1+1) algorithms. The derivation of the other values is detailed in App. E.
@@
Last but not least, Bet-And-Run algorithms, including 3-OneFifth, 3-Disc, and 2-DE/D-CMA, are robust, as proved in Sec. E.4 and demonstrated in the experimental results (Sec. 5 and App. I. However, they still require the choice of the embedded algorithms.

### v2

@@ 5.2 Discussion of experimental results
Robustness and statistical validation. Our experiments demonstrate that BBoxER achieves statistically significant performance gains without overfitting across a range of BBO algorithms (see P-value validation in App.˜J), these gains are also observed on Instruct models. Consistent with theory, the Bet-and-Run variant 3-OneFifth further improves performance and allows for larger budgets as shown in experiment 2. More discussion is provided in App.˜I.
@@
Computational cost. Our approach is memory-efficient, as it relies solely on LLM inference without the need to compute and store gradients, or large optimizer states. This allows us to fit a large batch size during training. The overhead introduced by the internal state of the BBO algorithm is negligible compared to the computational cost of LLM inference.
@@ 6 Conclusion
This paper introduced BBoxER  a principled post-training method that leverages comparison-based BBO algorithms for retrofitting. By compressing the training set into a comparison trace, measured in only a few bits, BBoxER offers strong theoretical guarantees on overfitting, robustness, and privacy. Despite this compression and the limited number of optimization iterations, BBoxER empirically demonstrates effective learning and transfer on GSM8k with contemporary LLMs, as well as empirical robustness to membership inference attacks. Theoretical insights are reinforced by experimental results, especially with Bet-and-Run algorithm variants. Importantly, this work is the first to bridge the LLM training landscape with the full BBO zoology. This connection, grounded in formal analysis, paves the way for safer, more robust, and adaptable language models, and opens the door to broader adoption of privacy-preserving training at scale, including completely private remote A/B testing or (post-)training directly on user satisfaction without any leakage of the underlying data.
Acknowledgments
JK thanks the Simons Foundation for support through the Collaborative Grant “The Physics of Learning and Neural Computation”.
Appendix Contents
@@ Appendix A Extended Related Work
Data poisoning alters training data to inject specific behaviors, either to degrade performance or to create targeted backdoors [70]. This includes subtle manipulations—e.g., poisoning just $0.001\%$ of training data in medical applications [71]—that remain undetected by conventional metrics. Attacks have proven transferable across quantized and full-precision models [41]. Recent work shows that inserting only a few percent of malicious preference pairs in RLHF [72] or DPO [73] data can dramatically shift model behavior [26, 25, 74, 75, 76, 77, 70], such as increasing output verbosity or sentiment polarity. These changes are activated by specific triggers and survive quantization. Theoretical work [54, 43] confirms that even small adversarial corruptions can be highly effective, especially in overparameterized models. However, certified robustness against poisoning in LLMs remains an open problem. Compared to e.g., [43], we provide an explicit algorithm with the desired stability assumptions.
@@
LLMs are exposed to risks of privacy and data extraction attacks [4, 78, 79, 42]. DP-SGD [45] introduces formal guarantees by adding Gaussian noise to clipped gradients, achieving $(\varepsilon,\delta)$-DP under strict accounting [80]. Nevertheless, it suffers from increased compute, memory use, and degraded utility. Specific implementations like Book Keeping [81], and extensions like adaptive clipping and layer-wise noise scaling aim to mitigate this [80]. PATE [47] aggregates teacher votes to train a student model privately. Parameter-efficient strategies like DP-BiTFiT [82], soft prompts [83], and selective DP [84] offer better trade-offs. Our method avoids the need for noisy gradient updates by enforcing a compression bottleneck and comparison-based optimization, achieving $(\varepsilon=0,\delta=0)$ for user input prompts and outputs in scenarious like A/B testing, by targeting high-fidelity memorization pathways directly. This aligns with broader insights that data retention precision is a root cause of privacy and misuse [85, 86]. Yu et al. [87] proposed a meta-framework for differentially private fine-tuning pretrained language models, achieving state-of-the-art privacy-utility tradeoffs on standard NLP tasks. Their approach was inspired by the success of parameter-efficient fine-tuning methods, allowing for simpler and sparser algorithms while maintaining high utility and providing rigorous privacy guarantees. In this case, by leveraging DP-SGD with careful gradient clipping and noise addition, the framework ensures that the influence of any single data point on the model’s parameters is bounded, thus protecting individual privacy. However, although the methods achieve strong performance, there is still a noticeable drop in accuracy compared to non-private fine-tuning, especially for smaller models or more complex tasks. Furthermore, the effectiveness of their DP fine-tuning approach is sensitive to hyperparameters such as learning rate, clipping norm, and noise multiplier, illustrating that careful fine-tuning is required to balance privacy and utility. Other works have focused on privacy-preserving strategies for instruction tuning and alignment of large language models. For instance, Yu et al. [88] introduced a two-stage framework that addresses privacy risks not only during training, but also in the instruction collection phase. In this case, they first fine-tune a pretrained LLM using differentially private optimization (DP-Adam) to generate synthetic user instructions, and then apply a DP histogram-based distribution matching technique to resample instructions such that they resemble the distribution of the original private data. This approach ensures that both annotators and the final alignment pipeline never see raw user inputs, thereby preventing both direct exposure and downstream memorization. Applied to LLaMA [89] and Phi-models [90], their approach achieved state-of-the-art performance in both supervised fine-tuning and RLHF settings, with utility comparable to non-private baselines. In particular, they highlight the importance of privacy-preserving data preparation in addition to model training.
@@
Reasoning tasks like GSM8k [56], SVAMP [62], MATH [57], and GSM-Symbolic [91] highlight the limitations of standard LLM (pre)training. Post-training approaches for reasoning tasks are now commonplace in LLM training. Enhancements via chain-of-thought prompting [92], complex reasoning extensions [93], and in-context learning [94, 95] are effective but often brittle. Reinforcement learning approaches such as GRPO [31] allow fine-tuning using unsupervised signals, opening the door for real-time user-driven improvement [96]. LoRA [97] and robust tuning [98] enhance adaptability but are prone to hallucination [99] and exploitative behaviors [100, 86]. Our approach combines low-rank adaptations with retrofitting techniques and avoids reward-based fine-tuning. It limits overfitting risks and prevents the amplification of memorized or poisoned content by emphasizing rank-based signal aggregation over token-level supervision. A somewhat complementary approach is self-improving [101]. In such a case, a large language model tries to improve its performance by self-fine-tuning with synthetic data generated by itself. Nevertheless, such approaches have mostly focused on performance, and they are not risk-free from data poisoning or leaking private data.
@@ Appendix B Examples of Modifiers
The BBoxER framework is described in Alg.˜1. This algorithm has various hyperparameters, such as a budget, a black-box optimization algorithm, and a $modified(m_{0},x)$ function which specifies how a vector $x$ proposed by the black-box optimization method is used to modify an initial model $m_{0}$. Here we present a few examples of such $modified(\dots)$ functions, involving broadcast, low-rank, or full updates.
@@ B.1 Examples of Modifiers for One Layer
In the present section $\times$ refers to pointwise multiplication and @@ refers to matrix multiplication.
@@
[…]
4: Return modelmodel
@@
Other update formulas include $0.01\times x$, $\times(Id+0.01\times x)$, and other constants than $0.01$. However, in most cases, we will use vectors $x$ rather than matrices, as explained in low rank and broadcast.
@@
When updating entire weight matrices, we prefer to work with low-rank (LoRA) updates (rank 1 in fact), defined as follows, when model.output_matrixmodel.output\_matrix has shape (hidden_dim,vocab_size)(hidden\_dim,vocab\_size) and $x$ has dimension hidden_dim+vocab_sizehidden\_dim+vocab\_size:
@@
[…]
5: Return modelmodel
@@
In some cases, we use broadcast rather than low-rank updates. $modified(m_{0},x)$ is then defined as follows, where $x$ as a row vector is repeated $M$ times to fit the shape of the matrix we want to update. In Sec.˜5 $x$ has dimension hidden_dimhidden\_dim and M=hidden_dimM=hidden\_dim:
@@
[…]
5: Return modelmodel
In App.˜I, $x$ has dimension vocab_sizevocab\_size when broadcast updates are applied to the output layer.
@@ B.2 Modifying Several Layers
In the examples above, we consider updates of one layer only. However, this can be naturally extended to updates of $l$ layers. Then, $x$ is split into $x_{1},\dots,x_{l}$, and each $x_{i}$ is used for updating a specific layer, as in Sec.˜B.1.
In our experiments in Sec.˜5 we performed the following updates:
@@
For all these experiments the update constant is $0.01$. A few additional experiments in App.˜I use a different constant.
@@ C.2 Probably Approximately Correct (PAC) Learning: A Brief Overview
  • Sec.˜C.2.1 presents the Bonferroni correction: whereas “non-uniform” large deviation inequalities estimate the difference between an empirical error and a generalization error for a single predefined classifier (for example when testing a single model on a test set), the Bonferroni correction provides bounds applicable uniformly on a finite set of classifiers (e.g., for selecting, with a validation set, a classifier in a list of classifiers obtained on the training set).
  • Sec.˜C.2.2 generalizes the Bonferroni correction to infinite families of classifiers: VC-dimension bounds, and others, are the most well-known bounds. They are based on limiting the “capacity” of a space of function, even when it is infinite.
  • Sec.˜C.2.3 presents other bounds for infinite families of functions, taking into account the optimization algorithm used for selecting a classifier: the idea is that even if the space of functions is large, we can get bounds if the optimization algorithm satisfies some stability assumptions. This last part of the state-of-the-art is the closest to our work: we get non-trivial bounds on deep learning models, independently of the possibly huge number of parameters.
@@ C.2.1 Bounds on the Generalization Loss based on Bonferroni
Let us see the generalization of large deviation inequalities seen in Sec.˜C.1.
@@
Thm˜1 extends this to more sophisticated cases, with $N(\omega,a,b)=\pi_{i}k_{i}(\omega,a,b)>>b$ (in many cases, exponential in $b$ e.g. $N(\omega,a,b)\leq 2^{b}$ and $k_{i}\leq 2$) instead of $N(\omega,a,b)=b$. In spite of being larger, this bound, exponential in $b$, is tighter than bounds based on capacity.
@@ C.2.2 Bounds on the Generalization Loss based on Capacity: Statistical Learning Theory from the 90s
[…]
@@ Appendix D Black-box Optimization Algorithms
The BBoxER algorithmic framework, as illustrated in Figure˜2, embeds an algorithm from the family of block-box optimization algorithms, which we now describe in detail.
Figure 2: Schematic view of BBoxER Alg.˜1. Our approach compresses the data relative to the current model, leading to a list $c$ of $b$ bits (Eq. 3, $b$ is the total budget). The final model depends only on those $b$ bits, giving perfect control over the number of bits transferred from the dataset to the model. In BBoxER, budgets $b$ are typically between 60 and 1500 bits, i.e., very small compared to the dataset size. The “Retrofitting compressor” refers to the part of Alg.˜1 which outputs the list $c$ of outputs of the $Compare(\dots)$ function: $c$ represents a compression of the dataset $D$ conditional on the initial model. “Retrofitting modifier” is the part of Alg.˜1 which outputs the final model, given the list $c$ and the initial model.
@@ D.1 Evolutionary Algorithms
For the reasons given in Sec.˜3, only Evolutionary Algorithms (EAs) have been used in the present work, as they are (at least for those uses in the present document) comparison-based, and hence de facto comply with the finite branching factor condition, which is mandatory for the theoretical results of Sec.˜4 to hold. Sec.˜D.2 discusses this choice, and its possible relaxation, while the current section introduces EAs more thoroughly, and details all the comparison-based algorithms used in the experiments of Sec.˜5.
@@
The first adaptive ES uses the so-called “one-fifth success rule” [114, 112], established by theoretical analyses on simple linear and quadratic objective functions, on which the optimal behavior is reached for a success rate (offspring better than parent) of approximately $0.2$. This is available in Nevergrad as the (1+1)-ES with 1/5th rule, named in this paper OneFifth, whose pseudo-code is given in Alg.˜2-left.
@@
When handling a full population ($\mu>1$), both CMA-ES and D-CMA apply the adaptive Gaussian mutation to a linear recombination of the $\mu$ individuals. The historical CMA-ES used the mean (all weights of the linear recombination are $1/\mu$), whereas the state-of-the-art CMA-ES uses weights that decrease with the rank of each individual. In the latter case, more information is used, and the branching factor increases (see Sec.˜E.5).
The simplest ESs (or EAs) are the (1+1)-EAs: the single current individual is modified, and the best of both becomes the new current one. While the already-mentioned OneFifth was the first historical successful adaptive EA (see again Alg.˜2-Left), another family of (1+1)-EAs is directly derived from the historical GAs (Genetic Algorithms), that optimize bitstrings: the mutation typically flips each bit with a given probability. In the continuous case, the bitflip is turned into a “Gaussian flip”: each value is replaced by a value sampled from a standard normal mutation (${\cal N}(0,1)$ in Alg.˜2). Several variants of such “discrete-inspired” algorithms exist in Nevergrad, depending on the probability of applying this Gaussian flip to each variable. The Discrete algorithm exactly mimics the standard setting in GAs by adopting a fixed $1/d$ probability; the Lengler [116, 117] algorithm uses a gradually decreasing schedule. The Uniform-Disc [118] algorithm uses a probability that is uniformly sampled in ]0,1]]0,1], and the FastGA [119] uses an $\alpha/d$ probability where $\alpha$ is sampled from a power-law distribution.
@@
Table 4: Bound on the average branching factor for various algorithms (see Sec.˜E.4, Sec.˜E.5, Sec.˜E.6).
@@
OneFifth 22
Lengler 22
FastGA 22
Portfolio 22
@@
DE, PSO 22
DE, PSO with curr-to-best 33
CMA-ES/D-CMA 22
@@
These algorithms were chosen because of their finite branching factor: Tab.˜4 lists the branching factors of all cited algorithms. The branching factor is 2 for all (1+1) algorithms. The derivation of the other values is detailed in App.˜E.
@@
Last but not least, Bet-And-Run algorithms, including 3-OneFifth, 3-Disc, and 2-DE/D-CMA, are robust, as proved in Sec.˜E.4 and demonstrated in the experimental results (Sec.˜5 and App.˜I. However, they still require the choice of the embedded algorithms.